Reports 1-1 of 1 Clear search Modify search
AdV-TCS (Wavefront sensing)
lumaca, nardecchia - 15:07 Monday 28 November 2022 (57924) Print this report
HWS-DET beam check

The goal of the shift was to verify the alignment of the HWS-DET sled beam after the work performed on the OMC.

First of all, we replaced the HWS-DET with the CCD, to verify the beam shape and eventually the presence of some clipping.

The beam, shown in figure 1, is pretty good, but something like a fiber seemed to be along the beam path (visible in fig.1, going from the top center to the mid left of the image).

We acted on EDB_Har_M1 (see fig. 2 for EDB Hartman optical layout) in vertical and horizontal directions, to slightly move the ongoing beam, trying to step away from the fiber.

The final beam is shown in fig. 3.

At this point, we replaced the CCD with HWS-DET and we checked the illumination (fig. 4).

To check the alignment of the beam on the NI-TM center, after moving EDB_Har_M1, we used the NI CH actuator, whose curvature appears faster on HWS maps wrt RH’s one. 

So we switched OFF the NI CH at 10.24 UTC. Then, we started an HWS-DET measurement and, at 10.45 UTC, we switched ON the NI at high power.

Soon after the NI CH appeared on the HWS-DET maps, pretty well centered, indicating that the HWS DET beam itself is well centered on the NI TM (fig.5).

So we stopped the HWS DET acquisition at 11.08 UTC and we restored the nominal NI CH power.

 

As a further check of HWS beam alignment on NI TM, we should take another HWS-DET measurement switching ON the NI RH: to be performed in the next days.

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×